# Philipp Wolters Johannes Gilg Torben Teepe Fabian Herzog Felix Fent Gerhard Rigoll

### v1

@@ 1 Introduction
Figure 1: Compared to previous real-time perception models our SpaRC model family achieves state-of-the-art performance in accuracy and inference speed. The inference speed is measured with a single consumer-grade RTX3090 GPU on nuScenes val.
@@
  • SpaRC achieves state-of-the-art performance on the nuScenes benchmark ($+2.9$ NDS and $+2.6$ mAP). Moreover, our findings generalize to the long-range and adverse conditions on the new TruckScenes benchmark and match the LiDAR-based baseline.
@@ 3.1 Radar Point Encoder
[…]
@@
[…]
@@ 4.1 Datasets
[…]
@@
[…]
@@
[…]
@@
[…]
Table 4: 3D Object Detection on TruckScenes val and test sets. The detection range is evaluated up to 150 m (denotes the official baselines of the benchmark).
@@ 4.4 Main Results
TruckScenes Val. Our architecture generalizes well to the new domain and sensor setup of TruckScenes, achieving a competitive NDS of 37.4 on the validation set (cf. Tab. 4). With more adverse conditions and longer detection ranges of up to 150 meters, the adoption of radar becomes more important, as SpaRC doubles the mAP (+11.8) over the current vision-only state-of-the-art. We provide more detailed information about the ranges, splits, and conditions in the Appendix.
@@ 4.7 Limitations and Outlook
For now, we rely on pre-processed radar data in the form of point clouds. However, lower-level radar representations like the high-dimensional radar cube also pose a promising research direction, with new challenges to efficient computations [54, 12, 8, 15, 14]. In future work, we want to explore how SpaRC and its sparse modeling can handle dense and raw radar tensors.
@@ B Implementation Details
down stride [$\times$2, $\times$2, $\times$2, $\times$2]
@@
Table 7: Point Transformer Backbone Settings.
@@
SpaRC C+R 43.8 29.6 1.1 40.0 8.6 18.3 16.5 14.1 40.8 14.3 0.0 41.9
@@ C Additional Ablations
Per Class Analysis. For a holistic understanding of SpaRC’s performance across different object classes (cf. Tab. 9), we present per-class mAP scores in Tab. 9. The radar-derived metric measurements significantly enhance the detection capabilities of our approach, enabling robust identification of both small-scale objects such as vulnerable road users (pedestrians and cyclists), which traditionally pose challenges for camera-only methods, as well as highly reflective metallic objects, including cars and trucks. This demonstrates the complementary nature of the multi-modal fusion approach for general robust depth estimation.
@@
SpaRC C+R 40.9 / 47.0 33.8 / 42.9 26.9 / 37.7 22.5 / 35.4
@@
SpaRC C+R 57.4 (+2.6) 58.6 (+1.6) 57.5 (+2.4) 36.1 (+5.7)
@@
Methods mAP Weather Daytime Lighting Season
SpaRC 22.5 23.2 13.7 18.8 24.5 31.0 20.2 17.6 13.4 23.4 13.4 16.7 17.8 24.9 15.7

### v2

@@ 1 Introduction
Figure 1: Compared to previous real-time perception models our 0.83137 0.15294 0.72941S0.7098 0.17255 0.70588p0.58431 0.19608 0.67843a0.45882 0.21961 0.65098R0.33725 0.23922 0.62745C0.21176 0.26275 0.6\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: model family achieves state-of-the-art performance in accuracy and inference speed. The inference speed is measured with a single consumer-grade RTX3090 GPU on nuScenes val.
@@
  • SpaRC achieves state-of-the-art performance on the nuScenes benchmark (+2.9+2.9 NDS and +2.6+2.6 mAP). Moreover, our findings generalize to the long-range and adverse conditions on the new TruckScenes benchmark and match the LiDAR-based baseline.
@@ 3.1 Radar Point Encoder
0.83137 0.15294 0.72941S0.7098 0.17255 0.70588p0.58431 0.19608 0.67843a0.45882 0.21961 0.65098R0.33725 0.23922 0.62745C0.21176 0.26275 0.6\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: C+R R50 $256\times 704$ 62.0 54.5 0.496 0.269 0.403 0.177 0.181
@@
0.83137 0.15294 0.72941S0.7098 0.17255 0.70588p0.58431 0.19608 0.67843a0.45882 0.21961 0.65098R0.33725 0.23922 0.62745C0.21176 0.26275 0.6\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: C+R R101 $512\times 1408$ 64.4 57.1 0.484 0.264 0.308 0.175 0.178
@@ 4.1 Datasets
0.83137 0.15294 0.72941S0.7098 0.17255 0.70588p0.58431 0.19608 0.67843a0.45882 0.21961 0.65098R0.33725 0.23922 0.62745C0.21176 0.26275 0.6\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: C+R V2-99 67.1 60.0
@@
0.83137 0.15294 0.72941S0.7098 0.17255 0.70588p0.58431 0.19608 0.67843a0.45882 0.21961 0.65098R0.33725 0.23922 0.62745C0.21176 0.26275 0.6\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: C+R V2-99 63.1 0.901
@@
0.83137 0.15294 0.72941S0.7098 0.17255 0.70588p0.58431 0.19608 0.67843a0.45882 0.21961 0.65098R0.33725 0.23922 0.62745C0.21176 0.26275 0.6\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: C+R V2-99 $640\times 960$ val 35.4 22.5 0.798 0.449 0.476 0.613 0.248
@@
HyDRa [67] C+R V2-99 $928\times 1952$ test 22.4 12.8 0.725 0.544 0.744 1.180 0.388
0.83137 0.15294 0.72941S0.7098 0.17255 0.70588p0.58431 0.19608 0.67843a0.45882 0.21961 0.65098R0.33725 0.23922 0.62745C0.21176 0.26275 0.6\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: C+R V2-99 $928\times 1952$ test 37.4 27.2 0.759 0.413 0.411 0.814 0.227
0.83137 0.15294 0.72941S0.7098 0.17255 0.70588p0.58431 0.19608 0.67843a0.45882 0.21961 0.65098R0.33725 0.23922 0.62745C0.21176 0.26275 0.6\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: C+R Intern-Image L $928\times 1952$ test 42.0 33.2 0.717 0.393 0.355 0.784 0.201
Table 4: 3D Object Detection on TruckScenes val and test sets. The detection range is up to 150 m (denotes the official baselines).
@@ 4.4 Main Results
TruckScenes Val. Our architecture generalizes well to the new domain and sensor setup of TruckScenes, achieving a competitive NDS of 37.4 on the validation set (cf. Tab. 4). With more adverse conditions and longer detection ranges of up to 150 meters, the adoption of radar becomes more important, as SpaRC doubles the mAP (+11.8) over the current vision-only state-of-the-art. We provide more detailed information about the ranges, conditions, and classes in the Appendix.
@@ 4.7 Limitations and Outlook
For now, we rely on pre-processed radar data in the form of point clouds. However, lower-level radar representations like the high-dimensional radar cube also pose a promising research direction, with new challenges to efficient computations [52, 12, 8]. In future work, we want to explore how SpaRC and its sparse modeling can handle dense and raw radar tensors.
@@ B Implementation Details
down stride $\times$2, $\times$2, $\times$2, $\times$2]
@@
Table 7: Radar Backbone Settings.
@@
0.80784 0.15686 0.72549S0.65882 0.18431 0.69412p0.5098 0.20784 0.66275a0.36078 0.23529 0.63137R0.21176 0.26275 0.6C\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: C+R 43.8 29.6 1.1 40.0 8.6 18.3 16.5 14.1 40.8 14.3 0.0 41.9
@@
In the breakdown of Fig. 5, we show that our sparse approach introduces minimal computational overhead while achieving significant performance gains.
@@ C Additional Ablations
LSA Configuration. Our ablation study (Tab. 12) on LSA demonstrates that K=32 (of all 644 queries) provides optimal performance, suggesting that focused local spatial context is sufficient for 3D object detection.
@@
0.80784 0.15686 0.72549S0.65882 0.18431 0.69412p0.5098 0.20784 0.66275a0.36078 0.23529 0.63137R0.21176 0.26275 0.6C\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: C+R 40.9 / 47.0 33.8 / 42.9 26.9 / 37.7 22.5 / 35.4
@@
0.80784 0.15686 0.72549S0.65882 0.18431 0.69412p0.5098 0.20784 0.66275a0.36078 0.23529 0.63137R0.21176 0.26275 0.6C\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: C+R 57.4 (+2.6) 58.6 (+1.6) 57.5 (+2.4) 36.1 (+5.7)
@@
TopK 4 8 16 32 128 644
mAP 39.9 43.8 45.1 46.4 46.4 45.1
NDS 51.5 53.7 54.2 54.9 54.6 54.5
Table 12: Ablation of query context aggregation window in LSA.
Figure 5: Latency breakdown of SpaRC model variants.
Per Class Analysis. For a holistic understanding of SpaRC’s performance across different object classes (cf. Tab. 9), we present per-class mAP scores in Tab. 9. The radar-derived metric measurements significantly enhance the detection capabilities of our approach, enabling robust identification of both small-scale objects such as vulnerable road users (pedestrians +8.9 mAP, cyclists +5.2 MAP, motorcycles +11.8 mAP), which traditionally pose challenges for camera-only methods, as well as highly reflective metallic objects, including cars (+27.2 mAP) and trucks (+17.5 mAP). This demonstrates the complementary nature of the multi-modal fusion approach for general robust depth estimation.
Methods mAP Weather Daytime Lighting Season
0.83137 0.15294 0.72941S0.7098 0.17255 0.70588p0.58431 0.19608 0.67843a0.45882 0.21961 0.65098R0.33725 0.23922 0.62745C0.21176 0.26275 0.6\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: 22.5 23.2 13.7 18.8 24.5 31.0 20.2 17.6 13.4 23.4 13.4 16.7 17.8 24.9 15.7
@@
Robustness Analysis Unlike methods relying on exact geometric projections (e.g. occupancy-based CRN), SpaRC handles misalignment through learnable positional embeddings and soft attention mechanisms. Our SFF uses column-wise topK attention rather than direct point-to-pixel projection, while RAR dynamically adjusts feature interactions based on learned spatial relationships. This learning-based approach inherently adapts to both spatial and temporal misalignments. Fig. 6 demonstrates SpaRC’s superior robustness to calibration errors (adding Gaussian noise to radar points), significantly outperforming HyDRa [67] and CRN [24] across all noise levels.
@@ D Qualitative Comparison
Figure 6: Robustness of mAP under different levels of radar misalignment (gaussian noise with std. $\sigma$) on nuScenes val set.